# Leveraging Diverse Data Generation for Adaptable Zero-Shot Dialogue State Tracking

### v1

@@ 5 Experiment Setup
MWOZ 2.4 IC-DST Hu et al. (2022) 175B* 58.7 62.1 53.2 54.9 71.9 51.4
RefPyDST King and Flanigan (2023) 175B* 68.8 74.5 56.6 68.2 68.5 76.1
T5 3B 49.2 63.2 26.0 71.7 29.8 55.8
      $+$D0T 3B 51.5 (+2.3) 69.1 29.9 73.2 29.2 56.2
[…]
@@
IC-DST Hu et al. (2022) 175B* 57.0 60.0 46.7 57.3 71.4 49.4
RefPyDST King and Flanigan (2023) 175B* 64.7 70.9 51.2 65.6 67.1 69.2
@@
T5 3B 44.7 57.1 22.8 60.6 30.9 52.2
      $+$D0T 3B 46.3 (+1.6) 61.5 27.3 62.3 29.1 51.7
[…]
Table 2: DST results on MultiWOZ, sorted within each division by avg. JGA. Parentheses indicate the improvement to performance resulting from training data diversity. * denotes the conjectured parameter size of OpenAI Codex. $\dagger$ denotes use of QLoRA.
6 Results
Table 2 presents zero-shot DST results on MultiWOZ 2.4 and 2.1 measured through Joint Goal Accuracy (JGA) as the evaluation metric. The results showcase the effectiveness of the presented data generation approach to create diverse training resources. By training on D0T, the zero-shot Llama model improves by $+6.7$ JGA on MultiWOZ 2.4, making it competitive with the much larger SoTA model using OpenAI Codex King and Flanigan (2023). The T5 results are more modest but still improved by $+2.3$ JGA.
The performance of our models see a massive performance dropoff on the MultiWOZ 2.1 test set compared to MultiWOZ 2.4, such as the zero-shot Llama$+$D0T suffering $-9.7$ JGA. This may be unusual, as largest performance dropoff from any baseline model when comparing MultiWOZ 2.1 vs 2.4 evaluation is only $-4.1$. Unfortunately, the MultiWOZ 2.4 corrections to the test set were made quite recently at the time of writing, thus most previous work only evaluates on MultiWOZ 2.1. Although Llama$+$D0T performs about as well as zero-shot ChatGPT and the OpenAI Codex-based approach IC-DST on MultiWOZ 2.4, the work by Gupta et al. (2022) still holds state-of-the-art on the MultiWOZ 2.1 benchmark at an impressive $65.9$ JGA, using a T5-11B model and fine-tuning all model parameters using 32 TPU machines. A question for future work is to what extent these large models benefit from training on diverse data with all parameters trainable.
@@ 7 Conclusion
In this work, we introduced a novel approach for generating domain-diverse DST data using instruction-tuned large language models (LLMs) to enhance zero-shot dialogue state tracking (DST) performance. Our approach involved the creation of a synthetic dataset comprising over $1,000$ task domains, which represents a significant expansion in the diversity of available training data. By automating the generation and annotation of dialogues, we addressed the bottleneck of limited domain diversity in existing DST datasets like MultiWOZ and SGD. The empirical results demonstrate the efficacy of our method: training zero-shot models on diverse synthetic data led to substantial improvements in Joint Goal Accuracy (JGA) on standard benchmarks, particularly on the corrected MultiWOZ 2.4 labels.
@@
This work shows that leveraging LLMs for data generation not only alleviates the high costs and logistical challenges of traditional data collection but also enables the creation of quality, diverse datasets that can lead to more generalized DST models. This initial work lays the groundwork for future research to further refine and enhance the quality of automatically generated data. By continuing to improve the diversity and correctness of synthetic datasets, we anticipate even greater advancements in zero-shot DST performance, driving the development of more robust and adaptable dialogue systems.
@@ Appendix D Implementation Details
T5-3B
Raffel et al. (2020) is a 3 billion parameter encoder-decoder transformer model trained on a variety of sequence-to-sequence tasks such as summarization and translation. Full fine-tuning for DST was performed with a learning rate of $41e-$ 4, a weight decay of $35e-$ 3, and batch size $128$.
@@
is a 13 billion parameter decoder-only transformer model trained on a variety of long-form texts, then further trained on instruction data using the Reinforcement Learning from Human Feedback (RLHF) technique Ouyang et al. (2022). Due to the computational expense of its 13B parameter size, the model was quantized using QLoRA Dettmers et al. (2023), which uses 4-bit nf4 quantization, and freezes the base model parameters while only training the parameters of a Low-Rank Adapter (LoRA) Hu et al. (2021). Training used a learning rate of $52e-$ 5, and a batch size of $256$.

### v2

@@ 6 Results
MWOZ 2.4 IC-DST Hu et al. (2022b) 175B 58.7 62.1 53.2 54.9 71.9 51.4
ParsingDST Wu et al. (2023) 175B 64.7 65.6 46.8 67.7 80.6 62.6
RefPyDST King and Flanigan (2023) 175B 68.8 74.5 56.6 68.2 68.5 76.1
T5-QLoRA 11B 47.1 63.9 24.1 65.5 29.4 52.9
    +D0T 11B 55.7 (+8.6) 68.1 32.0 72.3 50.6 55.8
Llama2-QLoRA 13B 59.2 62.2 44.9 69.8 49.1 70.2
      +ICL 13B 62.0 (+2.8) 74.7 44.9 69.8 49.1 71.3
    +D0T 13B 65.9 (+6.7) 74.4 56.4 76.0 54.7 68.3
    +D0T +ICL 13B 68.6 (+9.4) 76.8 56.4 78.8 54.7 76.1
@@
IC-DST Hu et al. (2022b) 175B 57.0 60.0 46.7 57.3 71.4 49.4
ParsingDST Wu et al. (2023) 175B 63.4 65.0 46.8 67.0 80.3 62.8
RefPyDST King and Flanigan (2023) 175B 64.7 70.9 51.2 65.6 67.1 69.2
@@
T5-QLoRA 11B 42.6 55.7 20.8 60.7 27.2 48.7
    +D0T 11B 49.9 (+7.3) 61.1 27.6 64.3 46.9 49.7
Llama2-QLoRA 13B 51.8 55.4 38.8 59.0 44.8 61.2
      +ICL 13B 54.0 (+2.2) 63.8 38.8 59.0 44.8 63.5
    +D0T 13B 56.2 (+4.4) 63.1 43.8 64.7 48.8 60.8
    +D0T +ICL 13B 58.5 (+6.7) 66.6 43.8 67.2 48.8 66.5
Table 3: Zero-shot DST results on MultiWOZ (JGA). Parentheses indicate the difference in performance compared to the baseline within base model groups. +D0T indicates training on D0T in an initial stage of training. +ICL indicates use of in-context demonstrations.
Impact of Domain-Diverse Training
Table 3 presents the results of the zero-shot DST evaluation. Training on the domain-diverse synthetic dataset D0T results in substantial performance gains across all models. On MultiWOZ 2.4, T5 and Llama2 gain +$8.6$ and +$6.7$ average JGA respectively. Gains on MultiWOZ 2.1 are more moderate at +$7.3$ for T5 and +$4.4$ for Llama2, which is expected as noisy gold labels make improvements less observable.
Interestingly, our models benefit from the gold label corrections of MultiWOZ 2.4 more than previous approaches. Llama2 +D0T +ICL benefits the most of any model from the MultiWOZ 2.4 corrections, indicating that it is punished for a substantial amount of correct predictions on MultiWOZ 2.1.
Llama2 demonstrated far better performance than T5 for both baseline and +D0T settings. With the improvements from D0T training, our Llama2 models achieve performance that is competitive with approaches based on language models of much larger ($\approx 175$ billion) parameter counts such as ChatGPT3.5 Heck et al. (2023); Wu et al. (2023) and OpenAI Codex Hu et al. (2022b); King and Flanigan (2023), and our best Llama2 +D0T +ICL model is within $0.2$% of the current SoTA.
Impact of In-Context Demonstrations
Adding in-context demonstrations to slot descriptions results in a consistent $2$-$3$% performance gain for both +D0T and baseline Llama2 models. This is consistent with previous work that tests the impact of in-context demonstrations Gupta et al. (2022). Encouragingly, the performance benefits of +ICL and +D0T appear to stack, yielding a combined improvement of +$9.4$ average JGA on MultiWOZ 2.4.
Comparison of Domain-Diverse Data
To further verify the effectiveness of D0T as a domain-diverse training resource, we compare against the most domain-diverse existing dataset, Schema-Guided Dialogues (SGD) Rastogi et al. (2020). We train a Llama2 model using the entire SGD training split as a first training stage to replace D0T training, before fine-tuning on MultiWOZ in the second stage to make a direct comparison. As shown in Table 4, the model leveraging D0T training outperforms a model that utilizes SGD instead. This demonstrates the power of the massively increased domain diversity covered by D0T, despite it being a synthetic dataset created with no human intervention. This result also validates the effectiveness of our automatic generation pipeline since it can yield useful training resources while only incurring a small fraction of the time and cost compared to traditional data collection methods.
TD F avg. attr. hotel rest. taxi train
SGD 65.1 76.0 51.6 76.8 53.5 68.0
SGD 61.8 75.6 45.1 77.0 46.8 64.5
D0T 65.9 74.4 56.4 76.0 54.7 68.3
D0T 66.3 78.8 53.9 75.0 53.0 71.1
Table 4: Zero-shot DST results on MultiWOZ 2.4 (JGA), comparing the efficacy of D0T versus SGD as a domain-diverse resource for stage one training. Llama2 is used as a base model with QLoRA training. TD: Stage one training dataset. F: Checked if domains similar to MultiWOZ are filtered out before training.
One limitation of evaluating SGD as a domain-diverse training resource on the MultiWOZ benchmark is that SGD contains an approximate superset of the domains in MultiWOZ. Consequently, the ability of SGD to train a domain-generalizable DST model is not tested. To address this, we simulate the effectiveness of SGD to improve zero-shot performance for new domains by filtering out all training examples that belong to a domain analogous to those seen in MultiWOZ. Specifically, we filter out the Travel, Hotel, Restaurant, RideShare, and Trains domains and train another baseline model using this filtered datatset. As shown in Table 4, zero-shot performance is impacted by -$3.3$ average JGA as a result of this filtering. Although D0T can be trivially extended to new domains using our automatic data generation pipeline, we similarly test its capability for training models that generalize to new domains by training a model using a filtered version of D0T. Filtering is performed by manually reviewing all $1,003$ domains and excluding any that include attractions, hotels, restaurants, taxis, trains, or general travel planning as a primary theme. Model performance remains virtually identical (+$0.4$) regardless of whether D0T domains are filtered based on similarity to MultiWOZ domains, which is evidence that the benefits of training on D0T generalize to unseen domains.
Impact of Trainable Parameter Size
We investigate the interaction between the parameter efficient training technique QLoRA and domain-diverse training by evaluating a variant of our T5 model with full finetuning and without quantization (i.e. without QLoRA). Additionally, a 3 billion T5 base model is compared to evaluate the impact of model size. Results are presented in Table 5. Consistent with previous work, we find that increasing model size yields substantial performance improvements on zero-shot DST. Whereas the T5-3B benefits from training on D0T, we observe a slight performance loss when training T5-11B, likely due to catastrophic forgetting when training on noisy D0T labels. Although QLoRA appears to moderately harm performance when training the T5-11B baseline, the T5-11B-QLoRA model actually achieves the best overall performance when first trained on D0T, likely due to the ability of QLoRA to protect against catastrophic forgetting.
model avg. attr. hotel rest. taxi train
3B 49.2 63.2 26.0 71.7 29.8 55.8
    $+$D0T 51.5 69.1 29.9 73.2 29.2 56.2
11B 53.8 65.0 27.6 71.0 37.5 68.2
    +D0T 52.4 70.3 29.1 66.8 36.1 59.9
11B-QLoRA 47.1 63.9 24.1 65.5 29.4 52.9
    +D0T 55.7 68.1 32.0 72.3 50.6 55.8
Table 5: Zero-shot DST results on MultiWOZ 2.4 (JGA), comparing 3B, 11B, and 11B-QLoRA variants of the T5 base model. +D0T indicates training on D0T in an initial stage of training.
Analysis of Training Stages
The efficacy of D0T as a training dataset for zero-shot DST is further investigated by comparing the performance of the Llama2 model at the conclusion of each stage of training. Table 6 presents results on the MultiWOZ 2.4 benchmark for the stage one model trained only on D0T versus the stage two model additionally trained on MultiWOZ. As expected, the second stage of training is revealed to be crucial as the stage one model achieves only $23.6\%$ average JGA. This reflects the effect of training on noisy dialogue state labels produced by automatic generation, which humans judged to have a slot-value pair correctness rate of $82\%$ [^5] Note that JGA is a more punishing metric than the percent of correct slot-values. Taken together with the results in Table 4, this result suggests that the benefit provided by D0T is due to its diversity rather than its overall quality compared to existing data. Further refinements to the automatic data generation pipeline presented in Section 3 to generate more accurate state labels may yield additional performance gains. An error analysis of stage one and stage two models is provided in Appendix E.
Stage avg. attr. hotel rest. taxi train
1 23.6 26.7 11.4 39.7 13.9 26.9
2 65.9 74.4 56.4 76.0 54.7 68.3
Table 6: Zero-shot DST results on MultiWOZ 2.4 (JGA), comparing Llama2 with QLoRA after training only on D0T (Stage 1) versus after additionally training on MultiWOZ (Stage 2).
@@ 7 Conclusion
The costly nature of DST data collection has been a limiting factor for the domain diversity of existing datasets for years. By introducing the first automatic data generation method capable of creating new domains and slot definitions for DST, this work both reveals and alleviates a performance bottleneck caused by the limited domain coverage of existing DST data. Training on the synthetic, domain-diverse D0T dataset produces substantial performance gains (e.g. +$6.7\%$ average JGA) for zero-shot DST, and this performance gain is stable even when testing on domains with no similar analog in synthetic data. These results show the power of domain diversity for training zero-shot DST models, as it allows our models to achieve competitive or better performance to LLM-based DST approaches with over 13.5×13.5$\times$13.5 × the parameters.
@@
The success of our data generation approach also demonstrates the potential of LLM-based data generation to alleviate the high costs of traditional data collection. Our work marks a pioneering step in the creation of similar fully automatic data generation approaches. By continuing to improve the diversity and correctness of synthetic datasets, we anticipate even greater advancements in zero-shot DST performance, driving the development of more robust and adaptable dialogue systems. We look forward to future research and application development in task-oriented dialogue that builds upon our experimental insights and released models and data.
8 Limitations
Redundancy of Slot Types
Although our presented data generation method successfully produces useful training data for zero-shot DST, it is important to note that this method does not produce a set of slot definitions where each slot is semantically unique. Our method attempts to maintain some consistency in tracking slots by modelling when requested slots are filled by a value. However, apart from tracking requested slots, slot-value update labels are generated relatively independently and without the notion of a centralized slot schema. This results in some cases, particularly across different dialogues belonging to the same domain, where slot labels are created with similar semantic meanings but different surface forms for slot names and descriptions. For training a zero-shot DST model this limitation is not an issue, since zero-shot DST models are expected to adapt to any provided slot name and definition to identify the correct value from the dialogue. However, the issue of inconsistent slot naming and lack of a centralized slot schema prevents datasets generated with our method from being used directly for few-shot training or DST evaluation.
Noise in Silver State Labels
Since our data generation technique is fully automatic, it is expected that some noisy silver labels of dialogue state occur. The $82.0$% slot-value correctness rate judged by our human annotators is interpretable as about 1 in 5 noisy slot-values. The limitation of this noise is that our experimental estimates of the impact of training data domain diversity on zero-shot DST are almost certainly under-estimates, as models trained on D0T were trained to predict this noise. Ideally, a dataset of similar diversity to D0T but with gold dialogue state labels would be used in our experiments; however, no such dataset exists, which is one of the primary motivations of our work. Our work thus serves as an investigation into the relationship between training domain diversity and zero-shot DST performance, but not one that conclusively quantifies this relationship. Future work should aim to reduce the noise in automatically generated DST labels or find more cost-efficient traditional data collection methods in order to achieve better experimental accuracy for measuring the impact of training domain diversity and in order to train higher-quality models.
9 Ethical Considerations
Risks
of this work are minimal; one risk introduced is through the use of GPT models to generate dialogue data, since it is theoretically possible for language model generations to populate synthetic dialogues with personal information of real people gathered from their training data. We believe the risk of this is low; after manually reviewing hundreds of dialogues in our D0T data, we observe that most potentially sensitive information is generated by GPT in anonymized form (e.g. the phone number 555-$5$).
Languages
used in this work are restricted to English, since it was required for all the authors to understand model outputs during prompt development and error analysis. The methodology presented in this work fundamentally language-agnostic however, and can be adapted to new languages by translating prompts. Since D0T is generated with a fully automatic method, analogous datasets in new languages can be created easily after prompt translation.
@@ Appendix D Implementation Details
is a 13 billion parameter decoder-only transformer model trained on a variety of long-form texts, then further trained on instruction data using the Reinforcement Learning from Human Feedback (RLHF) technique Ouyang et al. (2022). Due to the computational expense of its 13B parameter size, the model was quantized using QLoRA Dettmers et al. (2023), which uses 4-bit nf4 quantization, and freezes the base model parameters while only training the parameters of a Low-Rank Adapter (LoRA) Hu et al. (2022a) of rank 32. Training used a learning rate of $52e-$ 5, and a batch size of $256$, with no dropout or weight decay.
T5-11B
Raffel et al. (2020) is a 11 billion parameter encoder-decoder transformer model trained on a variety of sequence-to-sequence tasks such as summarization and translation. The T5 1.1 variant was used, following Gupta et al. (2022). QLoRA training used a rank of 32, alpha of 64, with a learning rate of $21e-$ 2 and batch size of $256$, with no dropout or weight decay. Full fine-tuning used a learning rate of $31e-$ 3 with weight decay $35e-$ 3.
Appendix E Error Analysis
Error Definition Stage Stage
1 2
Agent Value Miss No value is outputted for the indicated slot, even though the information is present in the system’s turns. 13 13
No Preference Indications of no preference are inappropriately understood, either by failing to recognize when no preference is given or by incorrectly interpreting an indication of no preference from the dialogue. 13 9
Value Change The appropriate value for the indicated slot has been updated in the dialogue turn, but the predicted value remains as the original. 10 19
Halluc- ination A value is predicted for the indicated slot that does not exist in the dialogue. 9 5
Miss No value is outputted for the indicated slot, even though the information is present in the user’s turns. 7 13
Wrong Value Information in the dialogue is incorrectly attributed to the indicated slot. 6 8
Other Errors not explained by any of the other error patterns. 16 11
Correct The predicted value for the indicated slot is correct, but is missing from the gold annotations in MultiWOZ due to an annotation mistake. 26 22
Table 7: Error analysis on 100 randomly sampled erroneous outputs on MultiWOZ 2.4 of the best-performing finetuned Llama-13B-Chat model with QLoRA training (Stage 2) and the same model trained only on D0T (Stage 1), before fine-tuning on MultiWOZ.
The impact of diverse DST training data is further investigated by conducting an error analysis on 100 randomly sampled errors from the best-performing Llama2 +D0T +ICL model. The model was evaluated for both Stage 1 (D0T training only) and Stage 2 (subsequent training on MultiWOZ), and the results of the error analysis can be seen in Table 7. As expected, some of the errors made by these models are due to slot semantics specific to the MultiWOZ task that are difficult to encode in a single-sentence slot description. For example, the dontcare value (represented as any to the model) is a frequent source of errors, as the model consistently overpredicts it in the Hotel domain. Many errors also stem from a slot being filled with a wrong value that does indeed appear in the dialogue, but does not quite fit the specifics of the definition of the MultiWOZ slot. However, the majority of errors made are due to limitations in the training formulation using the synthetic dataset. For example, the dialogues generated by GPT-3.5 rarely include corrections or clarifications where slot value would change, resulting in consistent errors when the user speaker changes their mind or self-corrects in MultiWOZ. Also, the military time format used in MultiWOZ for time slots was a consistent source of hallucinations, as this format rarely or never appears in the synthetic D0T data. Finally, the models frequently missed slot values entirely, particularly when the value originated from the system travel agent speaker.